a random sample of size n 16 is drawn from a normally distributed population. the sample mean is 98 with known standard deviation of 5. Construct a 98% confidence interval for the population mean.


1
Expert's answer
2022-04-17T08:46:14-0400

"CI=\mu \plusmn z \frac{\sigma}{\sqrt{n}}=98 \plusmn 2.33\frac{5}{\sqrt{16}}=98 \plusmn 2.91"